Use the substitution method to solve the linear system given. Express
the answer as an ordered pair of the form (x,y)
x + y = 5
2x + y = 6

Answer by gotpork1(3)

Use the substitution method to solve the linear system given. Express
the answer as an ordered pair of the form (x,y)
eq.1: x + y = 5
eq.2: 2x + y = 6

Step 1: Isolate for x or y in eq.1
x + y = 5
x = 5 - y

Step 2: Substitute x = 5 - y into eq.2
2x + y = 6
2(5 - y) + y = 6
(expand and simplify)
10 - 2y + y = 6
-1y = -4
y = 4


Step 3: Substitute y = 4 into eq.1
x + y = 5
x + (4) = 5
x = 5 - 4
x = 1


Therefore the solution is (1,4)